Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate
Supporting direct labor $ 16 per direct labor-hour
Order processing $ 182 per order
Custom design processing $ 264 per custom design
Customer service $ 424 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard
Model
Custom
Design
Number of gliders 14 3
Number of orders 1 3
Number of custom designs 0 3
Direct labor-hours per glider 26.50 34.00
Selling price per glider $ 1,625 $ 2,310
Direct materials cost per glider $ 460 $ 566

The company’s direct labor rate is $18 per hour.

Required:

Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)
Answer #1
Revenue 29680 =(1625*14)+(2310*3)
Costs:
Direct materials 8138 =(460*14)+(566*3)
Direct labor 8514 =(26.5*18*14)+(34*18*3)
Supporting Direct labor 7568 =(26.5*16*14)+(34*16*3)
Order processing 728 =(1*182)+(3*182)
Custom design processing 792 =3*264
Customer service 424
Total costs 26164
Customer margin 3516